a travels from rest with acceleration 8 metre per second square.Find its velocity when it has covered a distance of 100metre
1
Expert's answer
2013-08-30T13:59:01-0400
S = (V^2 - Vo^2)/2a
Where a - acceleration
Vo - velocity in the start = 0 m/s
V - velocity in the end
S = 100 m - distance
So,
V = sqrt(2as) = sqrt(2*8*100) = 40 (m/s)